the electrical resistance of the element in a platinum resistance thermometer at 100C, 0C and room temperature 75.000n, 63.000n and 64.000n respectively . determine the room temperature


1
Expert's answer
2022-02-22T08:52:55-0500

Solution;

"R=R_r(1+\alpha(T-T_r)"

"75=63(1+\alpha(100-0))"

"\frac {75}{63}=1+100\alpha"

"\frac{4}{21}=100\alpha"

Therefore;

"\alpha=\frac{1}{525}"

Then using the room temperature resistance;

"64=63(1+\frac{1}{525}(T-0))"

"\frac{64}{63}=1+\frac{T}{525}"

T=8.33°C